A​ $240 000 mortgage is amortized over 20 years. If interest on the mortgage is​ 3.39% compounded​ semi-annually, calculate the size of monthly payments made at the end of each month.  

A. ​$1,378.38

B. ​$1,375.47

C. ​$1,700.00

D. ​$1,184.36
Answer #1

B. ​$1,375.47

Size of monthly payment =-pmt(rate,nper,pv,fv)
=1,375.47
Where,
rate = Monthly Interest rate = 0.002805
nper = Number of period = 20*12 = 240
pv = Present value of loan = 2,40,000
fv = Future Value of Loan = 0
Working:
Calculation of equivalent monthly interest rate:
(1+i)^n = (1+i)^n Where,
(1+i)^6 = (1+0.01695)^1 Semi annual interest rate = 3.39%/2
1+i = 1.01695 ^ (1/6) = 0.01695
1+i = 1.002805253
i = 0.002805